Calculate the pressure on a floor by (a) one point
Answer #1

here,

Part a:
mass on one l leg, m = 62/ 4 = 15.5 kg

surface area, A = 0.021 cm^2 = 2.1*10^-6 m^2

Pressure, p = Force / Area

Pressure, p = m * g / A

Pressure, p = 15.5 * 9.81 / (2.1*10^-6)

Pressure, p = 72407142.857 N/m^2

Part b:
mass of elephant , m = 1598 kg

surface area, A = 803 cm^2 = 0.0803 m^2

Pressure, p = Force / Area

Pressure, p = m * g / A

Pressure, p = 1598 * 9.81 / (0.0803)

Pressure, p = 195222.665 N/m^2